# W-HMR: Human Mesh Recovery in World Space with Weak-supervised Camera Calibration and Orientation Correction

### v3

@@ 3.5 Ablation Study
[…]
@@
Camera Calibration We select PyMAF[54] as our baseline and consistently use ResNet50 as the backbone throughout our ablation study, with 431 sparse vertices as the downsampling points. The training and evaluation settings are all kept the same. From the Table 6, it is evident that the most significant improvements stem from adding the vertex regressor and global information. Including the vertex regressor enhances the model’s ability to reconstruct body mesh details. However, global information not only contributes independently but also aids in the functioning of camera calibration. A comparison between "V" and "V+F" reveals that solely adding the focal length predictor only has a noticeable impact on vertex accuracy. However, comparing "V+G" and "V+F+G" demonstrates that the presence of global information leads to more substantial improvements when combined with the focal length predictor. This finding further supports our idea that global information is crucial for recovering body orientation, predicting focal lengths, and correcting distortions. In summary, with the camera calibration assistance, W-HMR achieves a decent reconstruction accuracy, which provides a solid foundation for human body recovery in world space.
Table 7: The ablation study on body orientation correction. "spec-like" refers to using a similar method like SPEC to predict global pose directly. The numbers in brackets are the evaluation results on SPEC-MTP. This is the reason why we don’t choose the spec-like method. Its impressive results are based on overfitting and sacrificing generalization.
@@
${}^{\dagger}$(spec-like) 68.2(185.0) 43.6(63.3) 76.4(208.7)
${}^{\dagger}$ 82.1(118.7) 46.1(66.6) 93.3(133.9)
Orientation Correction We select two methods for evaluation. The first method is similar to SPEC. SPEC concatenates features and the camera parameters obtained from CamCalib, then inputs them into a regressor to directly recover the human body in world coordinate. As demonstrated in Table 7, the spec-like method significantly outperforms previous methods. However, we attribute this performance to overfitting. The model relies on massive parameters to memorize the global body orientation in the training set of SPEC-SYN, subsequently achieving excellent results on the test set. This is supported by the inferior performance of spec-like methods on SPEC-MTP. This indicates such method is prone to overfitting, which can lead to the misconception that the model can capture environmental information. What’s worse is that this design destroys the model’s original learned knowledge, making the overall reconstruction less precise. In contrast, our method offers a simpler structure, requires less training data, and exhibits greater robustness. Furthermore, other models based on camera coordinate can easily adopt our approach to achieve global pose recovery.
Table 8: Ablation study results on SPEC-MTP. The weights are from SPIN and are not fine-tuned. ${}^{\dagger}$ indicates using orientation output by OrientCorrect of pre-trained W-HMR. $*{}^{*}$ indicates HMR with an adaptive OrientCorrect module.
@@
${}^{\dagger}$ 149.6 75.4 179.4
$*{}^{*}$ 133.9 75.4 157.9
In order to verify the portability of OrientCorrect, we do related experiments, and the results are displayed in Table 8. To make the experiments more convincing, we choose HMR, a widely adopted model, as the baseline. We design two porting methods. One is to directly use the pre-trained OrientCorrect from W-HMR to predict the body rotation in the world space and then combine it with the pose and shape output from HMR to generate the final mesh. It can be seen that it brings nearly 14% improvement. From Table 3, it can be seen that using the camera rotations output from SPEC’s CamCalib to correct brings only about 10% boost. Moreover, the most impressive is the second porting method. We utilize the features and body orientation output by HMR to replace the original input part of OrientCorrect. In other words, we train an OrientCorrect adapted to HMR. The effect is pronounced, and the reconstruction accuracy is again improved by one grade. To avoid misunderstandings, you may think OrientCorrect is not as good as CamCalib or SPEC because the results in Table 8 are not as good as those of HMR and SPEC in Table 3. This is since we are based on different baselines. We use the pre-trained weights from SPIN directly and do not do any tuning, so the final results should be worse.
Design Details To find the optimal design for our model, we conduct further ablation experiments, as shown in Table 9. Firstly, we test different downsampling methods. Evidently, using markers for downsampling leads to better learning of human body features. Here, "markers" refer to mocap markers. The improved performance might be because mocap markers are a better proxy for inferring human body information [56]. However, since using mocap markers conflicts with the vertex regressor, we only employ markers in the pure parameter regression.
Table 9: Ablation experiments regarding the design details of modules. The first comparison involves vertex selection for downsampling. The second comparison examines the impact of different information inputs. The third comparison evaluates different designs of the vert graphomer. The fourth comparison evaluates different combinations of regressors. Within each comparison, all settings remain consistent except for the variations in module design.
@@
Lastly, we test various combinations of the hybrid regression module. Both approaches show advantages, but it is apparent that the combination of "2 param$+$1 vert" has clear benefits regarding model size and inference efficiency. Therefore, we ultimately select this combination as our final regression module.
4 Discussion
Of course, W-HMR is still not a perfect model, and for this reason, we show the failure cases in Fig. 9. We can find that our model is still not able to achieve accurate reconstruction when facing extreme poses (yoga, parkour), which may be due to the lack of relevant training data or may be limited by the iterative method. Secondly, when the human limbs are close to the body, there is often through-molding, which requires an extra optimization-based method to improve it or a human model with a higher degree of freedom. Third, W-HMR seems helpless against some severely distorted images taken at extremely rare angles (e.g., the image in the third row from the left). After all, as a weakly supervised method, W-HMR is limited in its ability to predict focal length in the absence of rich annotations for training. It is also true that W-HMR is unable to correctly establish the relationship between the person and the ground for images that lack background references. These will be one of the directions worth researching in the future.
Another important direction for human mesh recovery in world space is to capture the global trajectory of 3D human bodies. W-HMR mainly solves the problems caused by camera position and rotation. Although W-HMR can recover the human body posture in world space, it cannot obtain the absolute spatial position of the human body. Absolute spatial position is essential for the practical application of motion capture. However, at present, most related research still needs to rely on additional equipment to assist. This will be an important future research direction for motion capture based on monocular images.
@@ 4 Discussion
5 Conclusion
We proposed the first model capable of reconstructing human bodies accurately and reasonably in both camera and world coordinates. To achieve this, we designed a weak-supervised camera calibration method. "Weak-supervised" solves the problem of insufficient focal length labels. "Camera calibration" solves the problem caused by image distortion. The focal length predictor can generate appropriate focal lengths based on body distortion information, eliminating the dependence on background information and focal length labels, which are more suitable for human recovery tasks. Additionally, we introduced a new orientation correction module that is simpler and more robust than previous methods. Moreover, we proved that it can be ported to other models and work. With a carefully designed training paradigm and model framework, we achieved stable training and impressive performance across multiple datasets in both coordinates. We hope that our work will contribute to the study of human mesh recovery in world space. The codes and related materials will be kept updated. Everyone is welcome to acquire them from our project page https://yw0208.github.io/w-hmr/.

### v4

@@ IV-E Ablation Study
TABLE VI: The ablation study on the core block of the W-HMR. The experiments below utilize ResNet50 [36] as the backbone. In this table, ”V” refers to the vertex regressor. ”F” indicates whether the focal length predictor and FULL2 2D supervision are implemented. ”G” signifies the addition of global information during input.
@@
Camera Calibration We select PyMAF[13] as our baseline and consistently use ResNet50 as the backbone throughout our ablation study, with 431 sparse vertices as the downsampling points. The training and evaluation settings are all kept the same. From the Tab. VI, it is evident that the most significant improvements stem from adding the vertex regressor and global information. Including the vertex regressor enhances the model’s ability to reconstruct body mesh details. Moreover, global information contributes independently and aids in camera calibration. A comparison between ”V” and ”V+F” reveals that solely adding the focal length predictor only has a noticeable impact on vertex accuracy. However, comparing ”V+G” and ”V+F+G” demonstrates that the presence of global information leads to more substantial improvements when combined with the focal length predictor. This finding further supports our idea that global information is crucial for recovering body orientation, predicting focal lengths, and correcting distortions. In summary, with the camera calibration assistance, W-HMR achieves a decent reconstruction accuracy, which provides a solid foundation for human body recovery in world space.
TABLE VII: The ablation study on body orientation correction. ”spec-like” refers to using a similar method like SPEC to predict global pose directly. The numbers in brackets are the evaluation results on SPEC-MTP. This is the reason why we do not choose the spec-like method. Its impressive results are based on overfitting and sacrificing generalization.
@@
Ours-P(spec-like) 68.2(185.0) 43.6(63.3) 76.4(208.7)
Ours-P 82.1(118.7) 46.1(66.6) 93.3(133.9)
Orientation Correction We select two methods for evaluation. The first method is similar to SPEC. SPEC concatenates features and the camera parameters obtained from CamCalib, then inputs them into a regressor to directly recover the human body in the world coordinate. As demonstrated in Tab. VII, the spec-like method significantly outperforms previous methods. However, we attribute this performance to overfitting. The model relies on massive parameters to memorize the global body orientation in the training set of SPEC-SYN, subsequently achieving excellent results on the test set. This is supported by the inferior performance of spec-like methods on SPEC-MTP. This indicates that such a method is prone to overfitting, leading to the misconception that the model can capture environmental information. Worse, this design destroys the model’s original learned knowledge, making the overall reconstruction less precise. In contrast, our method offers a more straightforward structure, requires less training data, and exhibits greater robustness. Furthermore, other models based on the camera coordinate can easily adopt our approach to achieve global pose recovery.
TABLE VIII: Ablation study results on SPEC-MTP. The weights are from SPIN and are not fine-tuned. indicates using orientation output by OrientCorrect of pre-trained W-HMR. indicates HMR with an adaptive OrientCorrect module.
@@
HMR 149.6 75.4 179.4
HMR 133.9 75.4 157.9
To verify the portability of OrientCorrect, we do related experiments, and the results are displayed in Tab. VIII. To make the experiments more convincing, we choose HMR, a widely adopted model, as the baseline. We design two porting methods. One is to directly use the pre-trained OrientCorrect from W-HMR to predict the body rotation in the world space and then combine it with the pose and shape output from HMR to generate the final human motion. It can be seen that it brings nearly 14% improvement. From Tab. III, it can be seen that using the camera rotations output from SPEC’s CamCalib to correct brings only about 10% boost.
Moreover, the most impressive is the second porting method. We utilize the features and body orientation output by HMR to replace the original input part of OrientCorrect. In other words, we train an OrientCorrect adapted to HMR. The effect is pronounced, and the reconstruction accuracy is again improved by one grade. To avoid misunderstandings, you may think OrientCorrect is not as good as CamCalib or SPEC because the results in Tab. VIII are not as good as those of HMR and SPEC in Tab. III. This is because we are based on different baselines. We use the pre-trained weights from SPIN directly and do not do any tuning, so the final results should be worse.
Design Details To find the optimal design for our model, we conduct further ablation experiments, as shown in Tab. IX. Firstly, we test different downsampling methods. Using markers for downsampling leads to better learning of human body features. Here, ”markers” refer to mocap markers. The improved performance might be because mocap markers are a better proxy for inferring human body information [62]. However, since using mocap markers conflicts with the vertex regressor, we only employ markers in the pure parameter regression.
TABLE IX: Ablation experiments regarding the design details of modules. The first comparison involves vertex selection for downsampling. The second comparison examines the impact of different information inputs. The third comparison evaluates different designs of the vert graphomer. The fourth comparison evaluates different combinations of regressors. Within each comparison, all settings remain consistent except for the variations in module design.
@@
Lastly, we test various combinations of the hybrid regression module. Both approaches show advantages, but it is apparent that the combination of ”2 param$+$1 vert” has clear benefits regarding model size and inference efficiency. Therefore, we ultimately select this combination as our final regression module.
@@
V Discussion
Difference It is not the first time that predicting the distance between a person and a camera has been proposed. To understand our innovation, we take [63] as an example to discuss the differences between our work. Different Methods: The neural network in our work directly predicts the person’s distance from the camera. But in [63], they first calculate values $k$ based on the sizes of the bounding box and image. Then, they predict a correction factor $\gamma$ through a neural network and get the absolute depth value using $k$ and $\gamma$. Different Theories: Our work mainly relies on the degree of distortion of humans in the images. The model tends to predict shorter distances when humans are more distorted. But [63] mainly relies on the human size in images. A more miniature human means this human stands farther away from the camera. Different Purposes: We use predicted distance to calculate focal lengths. Our ultimate goal is to achieve a perspective camera model so that we can make full use of distorted images. As for [63], they want to recover accurate relative positions for multi-person pose estimation.
Limitation Of course, W-HMR is still not a perfect model, and for this reason, we show the failure cases in Fig 10. Our model can still not achieve accurate reconstruction when facing extreme poses (yoga, parkour), which may be due to the lack of relevant training data or the iterative method’s limitations. Secondly, when the human limbs are close to the body, there is often through-molding, which requires an extra optimization-based method to improve it or a human model with a higher degree of freedom. Third, W-HMR seems helpless against some severely distorted images taken at extremely rare angles (e.g., the image in the third row from the left). After all, as a weakly supervised method, W-HMR is limited in its ability to predict focal length without rich annotations for training. It is also true that W-HMR cannot correctly establish the relationship between the person and the ground for images lacking background references. These will be one of the directions worth researching in the future.
Future Another vital direction for human motion recovery in world space is to capture the global trajectory of 3D human bodies. W-HMR mainly solves the problems caused by camera position and rotation. Although W-HMR can recover the human body posture in world space, it cannot obtain the absolute spatial position of the human body. Absolute spatial position is essential for the practical application of motion capture. At present, most related research still needs to rely on additional equipment to assist. Some methods that can predict 3D pose [63, 51, 64] or SMPL [65, 66] in the world coordinate from monocular images have been proposed but still far from practical. This will be an important future research direction for motion capture based on monocular images.
VI Conclusion
We proposed the first model capable of reconstructing human bodies accurately and reasonably in both camera and world coordinates. To achieve this, we designed a weak-supervised camera calibration method. ”Weak-Supervised” solves the problem of insufficient focal length labels. ”Camera calibration” solves the problem caused by image distortion. The focal length predictor can generate appropriate focal lengths based on body distortion information, eliminating the dependence on background information and focal length labels, which are more suitable for human recovery tasks.
Additionally, we introduced a new orientation correction module that is simpler and more robust than previous methods. Moreover, we proved it can be ported to other models and work. We achieved stable training and impressive performance across multiple datasets in both coordinates with a carefully designed training paradigm and model framework. We hope that our work will contribute to the study of human motion recovery in world space. The codes and related materials will be kept updated. Everyone is welcome to acquire them from our project page https://yw0208.github.io/w-hmr/.